# What Lies Beneath? Exploring the Impact of Underlying AI Model Updates in AI-Infused Systems

### v2

@@ 3.3. Experiment Interface
  • Comparing Models Across Trials: After rating the current model, participants compared the model used in the current trial with the one from the previous trial (see Figure 4). The interface displayed relevant trial-specific details such as the query photo, model latency (i.e., time taken to retrieve results), the number of search results, whether the participant found a match, and their perceived accuracy. Participants then rated the similarity or difference between the models on a four-point Likert scale: very different, somewhat different, somewhat similar, or very similar. The system automatically labeled the model used in the first trial as Model A. Based on the participants’ comparison responses, the system assigned models in the subsequent trials as either Model A or Model B.
@@
  • Selecting a Preferred Model: After all 8 trials were completed, participants were shown a summary of their trial categorizations (see Figure 5), displaying which trials were categorized as Model A or Model B. Participants then selected their final preferred model for future photo identification tasks and provided a justification for their choice.
@@
In order to distinguish them from the actual underlying models (which were unknown to the participants), we will refer to these as Model A-UP and Model B-UP (user-perceived).
3.4. Facial Recognition Models
Similar to CWPS (Mohanty et al., 2019b), we used facial recognition models from Microsoft Azure for this study: the oldest available model, recognition_01 (released in 2017), and the most recent model, recognition_04 (released in 2021). We will refer to these as the old model and new model, respectively. According to the developer website (Urban, 2023), the new model is supposed to be the most accurate available, designed to handle challenging cases like faces with masks or facial hair. However, its effectiveness in specialized domains, such as historical photo identification, remains unclear. Previous benchmarking of the old model on CWPS showed that while effective at retrieving correct matches, it also retrieved a large number of search results (mostly false positives), increasing the likelihood of incorrect matches (Mohanty et al., 2020).
To better understand how these models differ in the context of historical photo identification, we conducted a small-scale benchmarking (see details in Appendix  B). This revealed notable differences between the two models: the new model consistently outperformed the old model in terms of average precision (i.e., ranking correct matches higher) and recall (i.e., finding more correct matches). While the new model retrieved significantly fewer results overall (an average of 20 compared to roughly 378 with the older model), those results were more accurate, with higher confidence scores for correct matches. The old model, in contrast, produced a large number of results but with more false positives, making it relatively less precise in this domain.
These differences in key metrics—such as the number of results, the ranking of more similar-looking candidates, and improved handling of diversity—make the old model and new model compelling candidates for this study. However, while these measurable differences exist, an important question remains: can users actually perceive these differences in practice? Since participants were unaware of which model was in use during each trial, the study explores whether users can rely on visible cues—such as fewer but more accurate results and better-ranked matches—to distinguish between the two models.
3.4.1. Latency in retrieving results
Prior work has demonstrated that latency—the time taken by a system to retrieve result—can significantly shape user perceptions of algorithmic systems, where even slight delays reduce engagement and satisfaction in search and chatbot contexts, but in some cases, slower responses may foster reflection or higher perceived quality (Park et al., 2019; Teevan et al., 2013; Brutlag, 2009; Gnewuch et al., 2018). In this study, we introduced an artificial delay to explore whether users consider latency when trying to distinguish between model changes and how it may influence their perception of model accuracy. Participants were assigned to one of four conditions: (i) Old Model retrieves results faster than New Model, (ii) New Model retrieves results faster than Old Model, (iii) both models retrieve results with random latency, or (iv) both models retrieve results with identical latency. For slower conditions, we introduced a 12-second delay, while standard conditions took 3 to 5 seconds.
@@ 3.7. Measurement
We collected several interaction and behavioral data points from 252 participants, each completing 8 trials, which resulted in a total of 2016 search sessions (see Table 1). In total, participants made 10646 comparison decisions across trials and completed 1764 pairwise model comparisons (7 per user). The new model was (randomly) assigned to 1015 search sessions, while the old model was assigned to 1001 sessions.
@@ 3.8. Analysis
  • Paired t-tests: Used to compare precision, recall, and false positive rates (FPR) between the models to assess improvements in identification performance. We also applied t-tests to examine participants’ ability to detect model changes and to compare perceived accuracy between the models.
  • Logistic mixed-effects models (GLMM): Used to predict participants’ ability to detect a model change (a binary outcome) and to analyze their final model preference. Key predictors included differences in perceived accuracy, latency, and search results between consecutive trials. GLMM was also applied to predict comparison decisions (e.g., positive vs. negative vs. uncertain decisions), using factors such as the underlying model type and time taken for individual comparisons. We implemented GLMM using the $glmer$ function in R (Bates et al., 2014).
  • Linear mixed-effects models (LME): Used to predict participants’ perceived accuracy of the models (a continuous outcome) and various behavioral metrics, including time taken for making comparison decisions, overall time spent browsing search results, number of results checked, and total number of comparisons. Key predictors included participants’ interactions with the models, such as the number of search results, the types of comparison decisions made (e.g., facial match or replica), and the underlying model type. We implemented LME using the $lmer$ function in R (Bates et al., 2014).
@@ 3.9. Findings
Out of 1764 pairwise model comparisons across 252 participants, the overall accuracy in detecting whether models had changed was 48.87%, close to random guessing. When the model changed, participants identified the change 56.98% of the time, but when the model remained the same, they accurately detected this only 38.38%. This difference in detection rates was statistically significant, with participants being better at identifying model changes than noticing when the model remained the same ($t(1665.5)=-7.91$, $p<0.001$). These findings contradict H1, as participants generally struggled to distinguish changes between models, especially when no change occurred.
To gain further insights, we examined participant comments after each trial. Some participants misinterpreted changes in the model’s behavior as variability or adaptations within a single model, rather than recognizing distinct models. One participant remarked: “Same AI? They got the 2 same matches,” after seeing similar outputs (e.g., common matches) across different trials, indicating that they thought they were interacting with the same model despite differences in other results. Another participant noted: “This model seems to be matching photos with consideration for ethnic features,” indicating they perceived the shift in behavior as an inherent adaptation of the same model, rather than a change to a different system.
@@
On the other hand, differences in observable characteristics like latency and result count did not significantly affect participants’ ability to distinguish between models (latency: $\beta=0.06054$, $z=1.005$, $p=0.315$; result count: $\beta=0.03753$, $z=0.635$, $p=0.525$). These findings directly contradict H2, as participants primarily relied on their subjective assessment of model performance rather than observable characteristics such as latency or the number of search results when distinguishing models between trials.
@@
Across 2016 trials, participants rated the perceived accuracy of the models on a scale (see Figure 3) from -100 (highly inaccurate) to 100 (highly accurate). Participants were tasked with determining whether the model they used in each trial was the same as or different from the previous trial, but they were never informed whether they were interacting with the old model or the new model. As shown in Finding 3.9.1, participants struggled to distinguish whether the model changed between trials, meaning any potential conscious mapping to the actual underlying model (old or new) was effectively lost.
Despite this, participants consistently rated the new model as significantly more accurate (mean accuracy = 33.62) compared to the old model (mean accuracy = 18.32; $t(1981.5)=-6.10$, $p<0.001$), thus supporting H3. It is worth noting that neither model was perceived as highly accurate, with both models receiving moderate accuracy ratings on the scale (see Figure 6).
@@
Finally, when participants interacted with the new model, they were more likely to rate the model as having higher perceived accuracy ($\beta=14.23,p<0.001$), even though they were unaware of which underlying model they were using. This is in line with H3. Interestingly, latency did not have a significant impact on perceived accuracy ($p=0.73$), while a higher number of search results had a small but significant positive effect on perceived accuracy ($\beta=5.97,p<0.001$).
@@ 3.9.5. User preferences were significantly driven by perceived model accuracy
When comparing participants’ decisions to ground truth across all 2016 search sessions, we observed only a marginal improvement in recall with the new model (62.64%) compared to the old model (56.05%), as confirmed by a two-sample Welch t-test ($t(1274.1)=-2.99$, $p=0.0028$). However, there was no significant difference in precision or false positive rates between the models (see Table 2). These findings suggest that LABEL:H4 was not fully supported, as improvements in Human-AI team performance with the more accurate model were minimal.
@@
Participants also took 0.61 seconds longer per comparison with the new model ($\beta=0.61$, $t=3.72$, $p=2\times 10^{-4}$). Interestingly, the time spent per comparison was linked to decision type: longer times were associated with an increased likelihood of positive ($\beta=0.046$, $z=15.24$, $p<2\times 10^{-16}$) and uncertain decisions ($\beta=0.040$, $z=13.30$, $p<2\times 10^{-16}$), while shorter times were linked to negative decisions ($\beta=-0.099$, $z=-24.62$, $p<2\times 10^{-16}$). These findings support  H5, as the underlying model significantly influenced user behavior.
@@ 3.9.8. Perceived model changes had negligible impact on user behavior
Contrary to  H6, perceived model changes had a limited effect on user behavior. Participants who believed the model had switched did not significantly alter the number of comparisons made ($\beta=0.24$, $t=0.57$, $p=0.57$), the time spent per session ($\beta=-0.06$, $t=-0.015$, $p=0.99$), or the mean time spent per comparison ($\beta=-0.27$, $t=-1.04$, $p=0.30$). However, participants did scroll through 2 more results on average when they perceived a model switch ($\beta=2.04$, $t=1.98$, $p=0.048$), suggesting slight exploratory behavior.
3.10. Key Takeaways
Our findings revealed that despite significant performance differences between the two underlying models, participants were unable to accurately distinguish between them, contradicting H1. Although the newer model demonstrated higher precision and recall, participants did not achieve any performance gains, refuting H4. This lack of improvement suggests that participants did not adapt their strategies when interacting with the more accurate model. Had they been explicitly informed of the model changes, participants might have adjusted their behavior by focusing on fewer but more accurate results. For instance, they could have prioritized evaluating top-ranked candidates rather than exhaustively scrolling through large result sets, which would have aligned their approach with the strengths of the new model. Such an adjustment could have led to better task performance and more efficient decision-making.
Moreover, our analysis shows that participants relied heavily on their personal interactions with the system to form perceptions of accuracy, which then influenced how they distinguished between models. Rather than using objective indicators like latency or result count, users trusted their subjective experiences, highlighting the need to communicate system updates in ways that reflect how changes will affect user interactions, rather than simply presenting technical improvements. The failure of H6, where perceived model changes did not significantly alter user behavior, further underscores that participants may not understand how to adjust their strategies when models change. Raising awareness about model capabilities and providing explicit guidance on how to adapt could help users optimize their performance and make better use of updated system functionalities.
Interestingly, while we observed similar false positive rates and human-AI team performance across both models, the newer model did induce different user behaviors. Participants spent more time deliberating when making positive matches or uncertain decisions and were quicker to rule out negative matches. While this did not result in overall performance gains, it suggests that the new model may be encouraging more thoughtful decision-making. Despite minimal improvements in measurable outcomes, the new model appears to be pushing users toward more ”reflective” behaviors, which could be beneficial with further guidance or in longer-term use. Overall, this underscores the need for explicitly informing users about model updates, which could lead to more effective use of the new model’s capabilities and a clearer understanding of how these changes influence both behavior and perceptions.
@@ 4. Study 2: Comparing User Perceptions of Old and New AI Models in a Real-World Deployment
Study 1 revealed that when model updates were invisible, users struggled to detect changes and did not fully capitalize on the more accurate model’s potential, leading to only marginal improvements in human-AI team performance. In this study, we explore whether explicitly informing users of a new model would lead to more effective use of its outputs and how their preferences and perceptions would differ when directly comparing two models.
We investigated how users of Civil War Photo Sleuth (CWPS)—a platform where the older facial recognition model had been in use—perceived and compared a newly introduced AI model. These users, already familiar with the older model, were informed that a ”new model” was available, but were given no specific details about its characteristics or improvements. Unlike Study 1, where model updates were hidden from participants, this study allowed users to explicitly toggle between the old and new models in a real-world environment, giving them the agency to explore differences between the models on their own terms.
@@ 4.1. Study Setup
We recruited 10 CWPS users who had actively used the platform’s facial recognition-based search workflow within the last six months. The group comprised nine males and one female, with an average age of 46 (min = 19, max = 62). Each participant expressed a certain degree of familiarity with the platform’s facial recognition search feature. Participants investigated the identities of 10 photos of their choice, either by uploading new images or selecting from existing ones on the site.
Over two weeks, they were instructed to use both models and document their reflections in a diary (see template in Appendix), including which model they found more useful and why, along with a usefulness rating from -100 (not at all useful) to 100 (highly useful). A mid-point reflection was collected after five photo investigations, followed by a final reflection at the conclusion of the study to assess their overall preferences and perceptions of both models. This study was approved by the university’s IRB.
@@ 4.2. Analysis
We reviewed and categorized participants’ diary entries, initially identifying 28 distinct codes, which were later grouped into 7 main themes (”Preference”, ”Perceived Accuracy”, ”Result Counts”, ”Other performance factors”, ”Misc. visual factors”, ”Unrelated to model”, ”Behavior”) discussed in the findings section.

### v5

@@ 3.4. Interface
  4. (4)
    Comparing Models Across Trials: After rating the current model, participants compared the model used in the current trial with the one from the previous trial (see Figure 4). The interface displayed relevant trial-specific details such as the query photo, model latency (time taken to retrieve results), the number of search results, whether the participant found a match, and their perceived accuracy. Participants then rated the similarity or difference between the models on a four-point Likert scale: very different, somewhat different, somewhat similar, or very similar.
    The labeling of trials followed a consistent process based on participants’ responses. The model in the first trial was always labeled as Model A by default. For each subsequent trial, participants compared the model in the current trial with the one from the previous trial and indicated whether they believed it was the same model or a different model:
    • If participants indicated the model was different, the current trial was assigned the opposite label from the previous trial (e.g., if the previous trial was labeled Model A, the current trial was labeled Model B, and vice versa).
    • If participants indicated the model was the same, the current trial retained the same label as the previous trial (e.g., if the previous trial was Model A, the current trial also remained Model A).
    This process created a sequence of labels (Model A or Model B) that reflected participants’ perceptions of whether the model changed or remained consistent across trials. It is important to note that these labels (Model A and Model B) represent user-perceived models and do not correspond to the actual underlying model (old or new) used in each trial, which was randomized independently.
@@
  5. (5)
    Selecting a Preferred Model: After all 8 trials were completed, participants were shown a summary of their trial categorizations (see Figure 5). This step in the workflow was designed to help participants organize their thoughts about the trials they categorized as Model A or Model B. The interface displayed key trial-specific information, such as task duration, the number of search results, whether a match was found, and the participant’s accuracy rating for each trial. Based on this summary, participants selected their preferred model for future photo identification tasks and provided a justification for their choice.
@@
To distinguish the user-perceived model labels from the actual underlying models (the old and new models, which were unknown to participants), we will refer to Model A and Model B as Model A-UP and Model B-UP (user-perceived) for clarity.
@@ 3.7. Measurement
We collected several interaction and behavioral data points from 252 participants, each completing 8 trials, which resulted in a total of 2016 search sessions (see Table 1). In total, participants made 10646 comparison decisions across trials and completed 1764 pairwise model comparisons (7 per user). The new model was (randomly) assigned to 1015 search sessions, while the old model was assigned to 1001 sessions. The sample size of 252 participants was chosen to ensure sufficient data for robust statistical analysis while remaining feasible within the constraints of the study design.
@@
The selected metrics were chosen to comprehensively evaluate the impact of model changes on both user performance and perceptions. Precision, recall, and false positive rate (FPR) are standard metrics for assessing the effectiveness of facial recognition systems, providing objective measures of identification accuracy, completeness, and error rates. Perceived accuracy and difference metrics were included to capture users’ subjective evaluations and the observable cues they relied on when comparing models. Specifically, we focused on the number of search results and response latency, as these are typical observable cues in retrieval tasks and directly influence user perception and behavior (Oulasvirta et al., 2009; Brutlag, 2009). Additionally, we recorded participants’ comparison decisions and the time taken to make these decisions, as these metrics are particularly relevant in the context of this study, where users are comparing search results to identify matches. Together, these metrics provide a balanced view of both objective system performance and subjective user experience, enabling us to address our research questions comprehensively.
@@ 3.8. Analysis
  • Paired t-tests: Used to compare precision, recall, and false positive rates (FPR) between the models to assess improvements in identification performance. We also applied t-tests to examine participants’ ability to detect model changes and to compare perceived accuracy between the models. This was used for the following findings: Sections 3.9.13.9.3, and 3.9.6.
  • Logistic mixed-effects models (GLMM): Used to predict binary outcomes such as participants’ ability to detect a model change (a binary outcome) and to analyze their final model preference. Key predictors included differences in perceived accuracy, latency, and search results between consecutive trials. GLMM was also applied to predict comparison decisions (e.g., positive vs. negative vs. uncertain decisions), using factors such as the underlying model type and time taken for individual comparisons. The fixed-effect coefficients ($\beta$) represent the magnitude and direction of the relationship between predictors and the log odds of the binary outcome. Exponentiating $\beta$ provides the corresponding odds ratio, which indicates the multiplicative change in odds for a one-unit change in the predictor. For instance, a one-unit increase in perceived accuracy difference ($\beta=-0.89630$) reduces the odds of detecting a model switch by approximately 59.2%, corresponding to an odds ratio of 0.408 ($e^{-0.89630}\approx 0.408$). We implemented GLMM using the $glmer$ function in R (Bates et al., 2014). This was used for the following findings: Sections 3.9.2 and  3.9.5.
  • Linear mixed models (LMM): Used to predict continuous outcomes, such as participants’ perceived accuracy of the models and various behavioral metrics (e.g., time spent making decisions, number of results checked). Key predictors included participants’ interactions with the models, such as the number of search results, the types of comparison decisions (e.g., facial match or replica), and the underlying model type. The fixed-effect coefficients ($\beta$) represent the magnitude and direction of the relationship between predictors and the outcome (e.g., perceived accuracy). For instance, each additional facial match response ($\beta=23.54$) increases perceived accuracy by 23.54 units, while scrolling through more results ($\beta=-4.75$) decreases it by 4.75 units. We implemented LMM using the $lmer$ function in R (Bates et al., 2014). This was used for the following findings: Sections 3.9.43.9.7, and  3.9.8.
@@ 3.9. Findings
Out of 1764 pairwise model comparisons across 252 participants, the overall accuracy in detecting whether models had changed was 48.87%, close to random guessing. When the model changed, participants identified the change 56.98% of the time, but when the model remained the same, they accurately detected this only 38.38%. This difference in detection rates was statistically significant, with participants being better at identifying model changes than noticing when the model remained the same ($t(1665.5)=-7.91$, $p<0.001$). These findings contradict 3.1, as participants generally struggled to distinguish changes between models, especially when no change occurred.
Participants’ comments revealed a notable inconsistency in how they interpreted perceived changes or continuity in the model’s behavior. Some participants misinterpreted changes in the model’s behavior as inherent variability or adaptation within a single model, rather than recognizing distinct systems. For instance, one participant remarked: ”Same AI? They got the two same matches,” after observing similar outputs (e.g., common matches) across different trials, suggesting they assumed consistency even when the model had changed. Conversely, other participants attributed perceived shifts in behavior to a different model, even when the underlying model remained the same. For example, one participant noted: ”This model seems to be matching photos with consideration for ethnic features”, indicating that they perceived these shifts as stemming from a new system.
@@
On the other hand, differences in observable characteristics like latency and result count did not significantly affect participants’ ability to distinguish between models (latency: $\beta=0.06054$, $z=1.005$, $p=0.315$; result count: $\beta=0.03753$, $z=0.635$, $p=0.525$). These findings directly contradict 3.1, as participants primarily relied on their subjective assessment of model performance rather than observable characteristics such as latency or the number of search results when distinguishing models between trials.
@@
Across 2016 trials, participants rated the perceived accuracy of the models on a scale (see Figure 3) from -100 (highly inaccurate) to 100 (highly accurate). Participants were tasked with determining whether the model they used in each trial was the same as or different from the previous trial, but they were never informed whether they were interacting with the old model or the new model. As shown in Finding 3.9.1, participants struggled to distinguish whether the model changed between trials, meaning any potential conscious mapping to the actual underlying model (old or new) was effectively lost. These ratings were based solely on their interaction with the outputs in each trial, without any knowledge of the underlying model’s identity. During analysis, we retrospectively mapped trials to the actual models (old or new) used, enabling us to compare perceived accuracy across the two models.
Despite this, participants consistently rated the new model as significantly more accurate (mean accuracy = 33.62) compared to the old model (mean accuracy = 18.32; $t(1981.5)=-6.10$, $p<0.001$), thus supporting 3.1. It is also worth noting that neither model was perceived as highly accurate, with both models receiving moderate ratings on the scale (see Figure 6).
@@
Finally, when participants interacted with the new model, they were more likely to rate the model as having higher perceived accuracy ($\beta=14.23,p<0.001$), even though they were unaware of which underlying model they were using. This is in line with 3.1. Interestingly, latency did not have a significant impact on perceived accuracy ($p=0.73$), while a higher number of search results had a small but significant positive effect on perceived accuracy ($\beta=5.97,p<0.001$).
@@ 3.9.5. User preferences were significantly driven by perceived model accuracy
When comparing participants’ decisions to ground truth across all 2016 search sessions, we observed only a marginal improvement in recall with the new model (62.64%) compared to the old model (56.05%), as confirmed by a two-sample Welch t-test ($t(1274.1)=-2.99$, $p=0.0028$). However, there was no significant difference in precision or false positive rates between the models (see Table 2). These findings suggest that 3.1 was not fully supported, as improvements in Human-AI team performance with the more accurate model were minimal.
@@
Participants also took 0.61 seconds longer per comparison with the new model ($\beta=0.61$, $t=3.72$, $p=2\times 10^{-4}$). Interestingly, the time spent per comparison was linked to decision type: longer times were associated with an increased likelihood of positive ($\beta=0.046$, $z=15.24$, $p<2\times 10^{-16}$) and uncertain decisions ($\beta=0.040$, $z=13.30$, $p<2\times 10^{-16}$), while shorter times were linked to negative decisions ($\beta=-0.099$, $z=-24.62$, $p<2\times 10^{-16}$). These findings support  3.1, as the underlying model significantly influenced user behavior.
@@ 3.9.8. Perceived model changes had negligible impact on user behavior
Contrary to  3.1, perceived model changes had a limited effect on user behavior. Participants who believed the model had switched did not significantly alter the number of comparisons made ($\beta=0.24$, $t=0.57$, $p=0.57$), the time spent per session ($\beta=-0.06$, $t=-0.015$, $p=0.99$), or the mean time spent per comparison ($\beta=-0.27$, $t=-1.04$, $p=0.30$). However, participants did scroll through 2 more results on average when they perceived a model switch ($\beta=2.04$, $t=1.98$, $p=0.048$), suggesting slight exploratory behavior.
3.10. Summary of Findings
Despite significant performance differences between the two models, participants in the online experiment were unable to accurately distinguish between them, contradicting 3.1. This resulted in no measurable improvement in performance, even when interacting with the newer, more accurate model, and thus refuting 3.1. In order to assess the models, participants relied heavily on their perceptions of accuracy, formed through personal interactions with the system, rather than using objective metrics such as latency or result count. The newer model did, however, lead to some changes in behavior, with participants spending more time deliberating on positive or uncertain matches, supporting 3.1.
@@ 4. Study 2: Comparing User Perceptions of Old and New AI Models in a Real-World Deployment
Study 1 revealed that when model updates were fully invisible, users struggled to detect changes and did not fully capitalize on the more accurate model’s potential, leading to only marginal improvements in human-AI team performance. In this study, we explore whether explicitly informing users of a new model would lead to more effective use of its outputs and how their preferences and perceptions would differ when directly comparing two models.
We investigated how users of Civil War Photo Sleuth (CWPS)—a platform where the older facial recognition model had been in use—perceived and compared a newly introduced AI model. These users, already familiar with the older model, were informed that a ”new model” was available, but were given no specific details about its characteristics or improvements. Unlike Study 1, where model updates were hidden from participants, this study allowed users to explicitly toggle between the old and new models (same as Study 1) in a real-world environment, giving them the agency to explore differences between the models on their own terms.
@@ 4.1. Study Setup
We recruited 10 CWPS users by posting an advertisement on social media to solicit interest and collaborating with the CWPS team to filter active users who had engaged with the platform in the past six months. The group comprised nine males and one female, with an average age of 46 (min = 19, max = 62). Each participant expressed a certain degree of familiarity with the platform’s facial recognition search feature. Participants investigated the identities of 10 photos of their choice, either by uploading new images or selecting from existing ones on the site.
Over two weeks, they were instructed to use both models and document their reflections in a diary (see template in supplemental material), including which model they found more useful and why, along with a usefulness rating from -100 (not at all useful) to 100 (highly useful). A mid-point reflection was collected after five photo investigations, followed by a final reflection at the conclusion of the study to assess their overall preferences and perceptions of both models. This study was approved by the university’s IRB.
@@ 4.2. Analysis
We reviewed and categorized participants’ diary entries, focusing specifically on responses provided in designated answer fields within the diary template. These fields captured participants’ thoughts on the old and new models, their perceived strengths and weaknesses, and their reflections on similarities and differences between the models.
One author conducted an open coding process to identify recurring concepts and observations within the answer fields. Through this process, we identified 24 unique codes, such as ”Speed,” ”Photo Quality,” ”Result Accuracy,” ”Facial Features,” ”Unrelated Features,” ”Mixed Results,” and ”Prefer Old/New/Both”. To ensure the robustness and consistency of the analysis, two authors collaboratively reviewed these codes and grouped them into seven overarching themes: ”Preference,” ”Perceived Accuracy,” ”Result Counts,” ”Other Performance Factors,” ”Misc. Visual Factors,” ”Unrelated to Model,” and ”Behavior”. This iterative process involved deliberation and consensus to ensure that the themes accurately reflected the diverse dimensions of participants’ reflections.
Finally, we categorized participants’ responses according to these themes, and the findings are presented in the subsequent sections of the paper.